Dalco manufacturing estimate that its weekly profit, P, in hundreds of dollars, can be approximately by the formula P= -2X SQUARE + 4 X +3, where X is the number of units produced per week, in thousand.
a. How many units should the company produce per week to earn the maximum profit?
b. Find the maximum weekly profit?

Answer by greenestamps(13200)



(a) The maximum value of the expression ax^2+bx+c (where a is negative) is at x = -b/2a. For the expression in this problem, x = -4/(2(-2)) = 1.

Since the units are thousands, the maximum profit will be when the number of units produced per week is 1,000.

(b) The maximum profit, in hundreds of dollars, is the value of the expression when x = 1: -2+4+3 = 5 (hundred dollars)

Not much of a business.... produce 1000 items in a week to make a profit of $500....